Can someone help me get the x y and z values for this system of linear equations question?
2x+4y-z=7
2x-4y+2z=-6
x+4y+z=0

Answer by richwmiller(17219)

z = 2x+4y-7,
z = -x+2y-3,
z = -x-4y
x = 1, y = 1/2, z = -3